# Chen, Simchi-Levi, and Wang

### v2

@@
Experimenting on MDPs
[…]
Experimenting on Markov Decision Processes with Local Treatments
@@
Utilizing randomized experiments to evaluate the effect of short-term treatments on the short-term outcomes has been well understood and become the golden standard in industrial practice. However, as service systems become increasingly dynamical and personalized, much focus is shifting toward maximizing long-term cumulative outcomes, such as customer lifetime value, through lifetime exposure to interventions. To bridge this gap, we investigate the randomized experiments within dynamical systems modeled as Markov Decision Processes (MDPs). Our goal is to assess the impact of treatment and control policies on long-term cumulative rewards from relatively short-term observations. We first develop optimal inference techniques for assessing the effects of general treatment patterns. Furthermore, recognizing that many real-world treatments tend to be fine-grained and localized for practical efficiency and operational convenience, we then propose methods to harness this localized structure by sharing information on the non-targeted states. Our new estimator effectively overcomes the variance lower bound for general treatments while matching the more stringent lower bound incorporating the local treatment structure. Furthermore, our estimator can optimally achieve a linear reduction with the number of test arms for a major part of the variance. Finally, we explore scenarios with perfect knowledge of the control arm and design estimators that further improve inference efficiency.
@@
Experimentation, A/B Testing, Lifetime Value, Markov Decision Process, Local Treatment
@@ 1 Introduction
However, as service systems grow more complex and dynamic, the focus is increasingly shifting from immediate or short-term outcomes to long-term effects on customers or users who are continually exposed to new interventions. For example, online retailers may employ strategies such as regular coupon distribution or cash back, aiming to trade-off short-term profit reductions for potential long-term revenue gains (Simester et al. 2006). Similarly, user interface updates on web platforms tend to have continual exposure to customers after their launches, thus impacting user experience and engagement over extended periods (Huang et al. 2023, Maystre et al. 2023). These practical scenarios require estimating the impact of long-term treatments on long-term cumulative outcomes, such as customer lifetime value, using relatively short-term observations from randomized experiments.
@@
In this work, we investigate the randomized experiments on Markov Decision Processes (MDPs), which are widely used to model dynamic service systems over extended time horizons including customer relationships (Pfeifer and Carraway 2000), inventory control (Gong and Simchi-Levi 2023), two-sided platforms (Johari et al. 2022), mobile health (Zhang et al. 2023) and many others. Additionally, as modern systems become increasingly complex with numerous states, the practical challenges of implementing entirely new global policies intensify. Such policies could be overly aggressive, in terms of posing significant risks and requiring substantial infrastructure changes. Consequently, many treatments in practice are becoming more fine-grained and localized. This means that the new treatments may only apply on a small number of targeted states locally. For instance, in the customer relationships, the customer’s states are referring to her recent behaviors or other contextual factors. An online marketplace or service platform might offer a “welcome-back” coupon to the customer only when she has not engaged with the service for an extended period (e.g., Sephora’s Welcome Back Event Dooley 2020). Similarly, physical retail locations like pharmacies and markets may provide incentives like cash back to the customer only when she makes in-store purchases (e.g., Bed Bath & Beyond Mohammed 2015). Likewise, a typical intervention for audio stream recommender systems may involve modifying banner promotions alone, without altering other system components throughout the iterative recommendation process (e.g., Spotify Maystre et al. 2023). These localized treatments are strategic, usually aiming to optimize the lifetime value of customers by tailoring experiences and offers to individual needs and histories. Given the prevalence of local treatments in practice, our research goal is to explore the potential of leveraging local treatment structure to optimize data utilization for efficient experimentation on the long-term cumulative reward within an MDP environment.
1.1 An illustrative Example On Customer Lifetime Value
@@
Drawing from the classic Recency, Frequency, Monetary value (RFM) framework, it is often posited by companies that the probability a customer will purchase at the end of any period is solely a function of her recency—the number of periods since her last purchase. Pfeifer and Carraway (2000) model recency as the state space of an MDP and characterize customer lifetime value via the value function, which is the discounted sum of revenue collected in the long-run.The state transitions occur as follows: each purchase resets the customer’s recency to one; conversely, if no purchase is made, the customer’s recency increases by one period. When a customer’s recency is high (for example, $\geq$5), Pfeifer and Carraway (2000) aggregate these instances into a single state, denoted as $s^{5}$, and regard it as representing an inactive customer.
@@ 1.2 Overview of Main Results
Efficient Inference under a Fixed Policy. Instead of simultaneously evaluating both the treatment and control policies, we initially focus on presenting estimation methods for the long-term cumulative rewards of a single, specific policy, say, either the treatment or the control. The two principal approaches used in the current literature are model-based estimation and temporal difference (TD) learning. Model-based estimation involves directly estimating all parameters of the MDP under a fixed policy to subsequently compute long-term rewards. Conversely, TD learning is more straightforwards to fit the long-term rewards according to the Bellman equation. Despite their distinct approaches, we establish that both methods are equivalent and demonstrate asymptotic normality. This means that the two estimators yield identical results and share confidence intervals. This equivalence allows the theoretical results derived from one method to be applicable to the other, providing the flexibility to select the method that is simpler to analyze. Moreover, we derive a variance lower bound for all unbiased estimators using a constrained Cramér-Rao bound. This highlights the inherent challenges of the inference task and confirms that both methods are asymptotically efficient compared to all other unbiased estimators, which may be of independent interest for the reinforcement learning (RL) literature.
Variance Reduction Utilizing the Local Treatment Structure. To accurately infer the ATE, we simultaneously estimate the long-term cumulative rewards for both the treatment and control policies. Particularly, we follow the spirit of A/B testing and apply the treatment and control policies randomly at each time period with equal probability. Significant challenges arise from the fact that the data collected are generated by a mixed policy—neither solely the treatment nor the control policy—and our focus is on the long-term cumulative rewards, not just instant rewards. We first establish an unbiased estimator for the ATE, treating the treatment and control policies as completely distinct, without leveraging the SST structure. This estimator is shown to be asymptotically normal, and its variance is almost optimal when the SST structure is not used, as confirmed by our derived variance lower bound.
@@
More importantly, we introduce an estimator that incorporates the SST structure by sharing data collected from states where the new treatment is not applied. This estimator can always achieve a smaller variance, demonstrating the benefits of utilizing the SST structure. Importantly, when expanding to multiple test arms, we illustrate that a major part of the asymptotic variance can be reduced linearly with the number of test arms compared to the estimator that does not use the SST structure. Additionally, we derive another variance lower bound for all unbiased estimators even with the knowledge of the SST, validating the optimality of our proposed estimator.
Further Variance Reduction with Perfect Information about the Control Arm. In many applications, the control policy has been operational as the default setting within systems for an extended period, yielding a wealth of offline data. Inspired by this abundance of historical information, we explore a scenario where we possess perfect knowledge of the control policy, including its state transition probabilities and immediate rewards. In response, we introduce a new estimator that is asymptotically normal, leveraging both our comprehensive understanding of the control policy and the SST structure. Our findings demonstrate that this approach can further reduce the asymptotic variance, enhancing the estimator’s effectiveness.
@@
Having demonstrated the variance reduction achieved through SST, we also expand our framework and generalize the results to encompass more general local treatments. This extension implies that the new treatment may be applicable to multiple states, broadening the scope and applicability of our results. We also test the effectiveness of our design on a real-world Sepsis simulator. Moreover, we reveal that under SST, the treatment effect remains consistently positive (or negative) for all states before and after the intervention. This indicates that if the new treatment benefits one state, it benefits all others as well. Also, we discuss that our information sharing concept is orthogonal to and compatible with the bias-variance trade-offs (e.g., Farias et al. 2022) and function approximation approaches (e.g., Shi et al. 2020, Syrgkanis and Zhan 2023) developed for the global treatments. This means that many developments under the global treatments can be further improved with the presence of the local treatments by our information sharing idea.
@@ 1.3 Related Works
Structure-aware Experiment Design. Experimental design originated in economics with basically model-free methods. In recent years, researchers in operations research and management science instead realize that there are many useful structures in operations scenarios to understand, analyze and improve experimentation. Bojinov et al. (2023), Hu and Wager (2022), Ni et al. (2023), Jia et al. (2023), Xiong et al. (2024) consider the switchback design under various scenarios, from temporal interference to spatial interference, from stationary environment to non-stationary environment, and from theory driven perspective to data driven perspective. Glynn et al. (2020), Munro et al. (2024), Wager and Kuang (2021), Bajari et al. (2021), Johari et al. (2022), Zhu et al. (2024), Weng et al. (2024) study the experiments happening on two-sided markets and a common practice is to model the system as a continuous-time Markov chain. Li et al. (2015) leverage the sparsity structure of cross-product demand elasticities to greatly reduce the number of experiments required. Li et al. (2023) carefully take the queuing structure into account and propose efficient estimators under the stochastic congestion. Xiong et al. (2023) investigate multi-unit settings with varying treatment starting times. Zhan et al. (2021, 2023) study the post-experiment inference for contextual bandits, which does not capture the state transition as our MDPs. Simester et al. (2020) pioneer at considering treatment structures and suggest comparing outcomes only for customers exposed to different treatment and control actions to improve efficiency empirically. Compared to theirs, we further incorporate complex transition dynamics and provide theoretical guarantees of estimations. The long-term effect with the long-term treatment under the short-term measurement is recently studied by Huang et al. (2023). This paradigm of structure-aware experiment design has witnessed great success in theory and practice in above works with either significant reduction of bias or substantial improvement of efficiency.
@@
Experimenting on MDPs with General Treatments. Experimenting on MDPs is more relevant to our work. Many works adopt MDPs to capture the dynamical system and investigate the long-term outcomes (Pfeifer and Carraway 2000, Murphy 2003, Simester et al. 2006, Tran et al. 2023, Shi et al. 2020, 2023, Farias et al. 2022, 2023), where most previous works consider general treatment, i.e. a treatment will be posed to every state of MDP. The seminal work by Farias et al. (2022) studies the average reward setting. They prove the first asymptotic lower bound for unbiased inference on MDPs and show that this lower bound can sometimes grow exponentially with the number of states. They also construct an LSTD(0)-type OPE estimator that is efficient and achieves this lower bound. In our work, we show that similar phenomena can also be observed in the discounted reward settings. Moreover, Farias et al. (2022) propose an alternative estimator that achieves a notable bias-variance trade-off with an upper bound of asymptotic variance. In contrast, all estimators discussed in this paper are asymptotically unbiased with exact formulae of asymptotic variance preferable for accurately determining the local treatment effect. We instead manage to reduce variance by taking into account the treatment structure. We also discuss the adaptation of their bias-variance trade-off estimator to the discounted reward setting (Farias et al. 2023) in Section 9.2. Furthermore, given the variance lower bound, some recent work like Shi et al. (2023), Zhang et al. (2023) considers asymptotic results with linear function approximation. Note that our methods are orthogonal to the bias-variance trade-off estimator and the function approximation method, and thus are compatible with them.
@@ 1.4 Structures and Notations
In Section 2, we formally introduce the model and the formulation, introducing the MDP, the treatment policy and the control policy, and the average treatment effect that we care about. Section 3 presents the estimation methods for the long-term cumulative rewards of a single, specific policy. In Section 4, we discuss the idea of classical A/B testing in our setting and how to estimate the long-term cumulative rewards for the treatment and control policies from A/B testing. Section 5 provides a new estimator based on the idea of information sharing for the local treatment. Section 6 studies the setting with perfect information of control. Section 7 generalizes the analysis to the general local treatments. In Section 8, we provide some numerical results verifying the idea of information sharing. Section 9 covers the discussion on some useful properties of the local treatment, and the compatibility with other popular techniques in RL. We conclude our paper in Section 10.
@@ 2 Model
where $10\leq\gamma<$ 1 is a discount factor that reflects the trade-off between instant and future rewards. The value function has been widely applied to quantify long-term outcomes in various fields including marketing (Pfeifer and Carraway 2000, Simester et al. 2006), ride-hailing platforms (Shi et al. 2023), and video streaming service (Tran et al. 2023, Maystre et al. 2023). We further use ${\bm{V}}^{\pi}\in\mathbb{R}^{K}\coloneqq(V^{\pi}(s^{1}),\cdots,V^{\pi}(s^{K}))^{\top}$ to denote value functions of all states. Moreover, associated with each $\pi$, ${\bm{r}}^{\pi}\in\mathbb{R}^{K}$ is the expected reward under each state under policy $\pi$, defined as ${\bm{r}}^{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}r(s,a)\pi(a\mid s)$ for $s\in{\mathcal{S}}$. ${\bm{P}}^{\pi}\in\mathbb{R}^{K\times K}$ is the transition matrix under $\pi$ with ${\bm{P}}^{\pi}(s,s^{\prime})\coloneqq\sum_{a\in{\mathcal{A}}}P(s^{\prime}\mid s,a)\pi(a\mid s)$ for $s,s^{\prime}\in{\mathcal{S}}$.
We then introduce a special class of treatment policies that we call the Single-State Treatment (SST). Specifically, there exists a pre-determined “crucial” state, which without loss of generality, we assume to be $s^{1}$. The treatment of interest is applied exclusively when the system’s state, $s_{t}$, coincides with $s^{1}$. Formally, we assume that ${\mathcal{A}}=\{{\mathfrak{c}},{\mathfrak{t}}\}$ regardless of states, while the action ${\mathfrak{t}}$, representing the new treatment, is only applicable in the crucial state $s^{1}$, while ${\mathfrak{c}}$ denotes the control or default action. This restriction ensures that the treatment is targeted and its effects are isolated to the crucial state.
@@
Before we conclude the session, we would like to introduce an important measure, which plays an important role in our theoretical results and is also commonly used in the RL literature (Agarwal et al. 2022).
Definition 2.1 (Discounted State-Visitation Measure)
Given a policy $\pi$, for any two states $s,s^{\prime}$, the Discounted State-Visitation Measure is the infinite discounted summation of the probability of visiting state $s$ starting from state $s^{\prime}$, i.e.,
$\rho^{\pi}_{s^{\prime}}(s)$ $\coloneqq\sum_{i=0}^{\infty}\gamma^{i}\mathbb{P}^{\pi}(s_{i}=s\mid s_{0}=s^{\prime})\,.$
Intuitively, the discounted state-visitation measure quantifies the likelihood that policy $\pi$ will lead to state $s$ in the future, starting from the initial state $s^{\prime}$. When the discounted factor $\gamma$ approaches $1$, this measure closely approximates the expected number of times the state $s$ is visited. This metric effectively captures the underlying dynamics of the MDP. Given that our intervention is exclusively applied to state $s^{1}$, the frequency with which state $s^{1}$ is visited becomes a crucial factor in assessing the impact of the treatment.
3 Preliminaries: Inference under A Fixed Policy
Before presenting the methods for experimenting with and comparing different policies, in this section, we first discuss how to perform inference on the value function under a fixed policy $\pi$, as an important cornerstone. Estimating the value function within MDPs is not a new task in the literature. There are two streams of methods, both of which are reliant on the well-known Bellman Equation, i.e.,

### v3

@@
Experimenting with Treatment Locality
@@
Utilizing randomized experiments to evaluate the effect of short-term treatments on the short-term outcomes has been well understood and become the golden standard in industrial practice. However, as service systems become increasingly dynamical and personalized, much focus is shifting toward maximizing long-term outcomes, such as customer lifetime value, through lifetime exposure to interventions. Our goal is to assess the impact of treatment and control policies on long-term outcomes from relatively short-term observations, such as those generated by A/B testing. A key managerial observation is that many practical treatments are local, affecting only targeted states while leaving other parts of the policy unchanged. This paper rigorously investigates whether and how such locality can be exploited to improve estimation of long-term effects in Markov Decision Processes (MDPs), a fundamental model of dynamic systems. We first develop optimal inference techniques for general A/B testing in MDPs and establish corresponding efficiency bounds. We then propose methods to harness the localized structure by sharing information on the non-targeted states. Our new estimator can achieve a linear reduction with the number of test arms for a major part of the variance without sacrificing unbiasedness. It also matches a tighter variance lower bound that accounts for locality. Furthermore, we extend our framework to a broad class of differentiable estimators, which encompasses many widely used approaches in practice. We show that all such estimators can benefit from variance reduction through information sharing without increasing their bias. Together, these results provide both theoretical foundations and practical tools for conducting efficient experiments in dynamic service systems with local treatments.
@@
Customer Lifetime Value, Randomized Experiments, Local Treatment
@@ 1 Introduction
However, as service systems grow more complex and dynamic, the focus is increasingly shifting from immediate or short-term outcomes to long-term effects on customers or users who are continually exposed to new interventions. For example, online retailers may employ strategies such as regular coupon distribution or cash back, aiming to trade-off short-term profit reductions for potential life-time revenue gains (Simester et al. 2006). Similarly, user interface updates on web platforms tend to have continual exposure to customers after their launches, thus impacting user experience and engagement over extended periods (Huang et al. 2023, Maystre et al. 2023). These practical scenarios require estimating the impact of long-term treatments on long-term outcomes, such as customer lifetime value, using relatively short-term observations from randomized experiments.
@@
In this work, we investigate the randomized experiments on Markov Decision Processes (MDPs), which are widely used to model dynamic service systems over extended time horizons including customer relationships (Pfeifer and Carraway 2000), inventory control (Gong and Simchi-Levi 2023), two-sided platforms (Johari et al. 2022), mobile health (Zhang et al. 2023) and many others. Additionally, as modern systems become increasingly complex with numerous states, the practical challenges of implementing entirely new global policies intensify. Such policies could be overly aggressive, in terms of posing significant risks and requiring substantial infrastructure changes. Consequently, many treatments in practice are becoming more fine-grained and localized. This means that the new treatments may only apply on a small number of targeted states locally. For instance, in the customer relationships, the customer’s states are referring to her recent behaviors or other contextual factors. An online marketplace or service platform might offer a “welcome-back” coupon to the customer only when she has not engaged with the service for an extended period (e.g., Sephora’s Welcome Back Event Dooley 2020). If she is a daily active user, she might not receive it. Similarly, physical retail locations like pharmacies and markets may provide incentives like cash back to the customer only when she makes in-store purchases (e.g., Bed Bath & Beyond Mohammed 2015). Likewise, a typical intervention for recommender systems may involve precise exposure, only taking effect when user click into certain pages such as banners or author profiles, without altering other system components throughout the iterative recommendation process (e.g., Spotify Maystre et al. 2023 and Douyin). These localized treatments are strategic, usually aiming to optimize the lifetime value of customers by tailoring experiences and offers to individual needs and histories. Given the prevalence of local treatments in practice, our research goal is to explore the potential of leveraging treatment locality to optimize data utilization for efficient A/B testing on long-term effects within an MDP environment.
1.1 An Illustrative Example: Customer Lifetime Value
@@
Drawing from the classic Recency, Frequency, Monetary value (RFM) framework, it is often posited by companies that the probability a customer will purchase at the end of any period is solely a function of her recency—the number of periods since her last purchase. Pfeifer and Carraway (2000) model recency as the state space of an MDP and characterize customer lifetime value via the value function, which is the discounted sum of revenue collected in the long-run.
@@ 1.2 Overview of Main Results
First, we design an efficient estimator of ATE in general A/B testing on MDPs, where the treatment does not need to be local. Specifically, we show that the two principal approaches from the reinforcement learning (RL) literature, the model-based estimation and temporal difference (TD) learning, remain effective for estimating ATE if the data are split correctly. Our proposed estimator is asymptotically normal and unbiased. Moreover, we show that no unbiased estimator can achieve a smaller variance without considering treatment locality. This result builds on a new efficiency proof framework via the constrained Cramér-Rao bound.
@@
We then introduce an estimator that incorporates the SST structure by sharing data collected from states where the new treatment is not applied. While data sharing naturally increases the sample size for both the treatment and control policies, it also introduces additional correlations that complicate the overall variance. Fortunately, we are able to show that this estimator can always achieve a smaller variance compared to the original estimator that does not share data. Our analysis further reveals that the variance reduction arises specifically from the component of the variance associated with the states that share data. Importantly, when extending to multiple test arms, we illustrate that a major part of the asymptotic variance can be reduced linearly with the number of test arms compared to the estimator that does not use the SST structure. Additionally, we derive another variance lower bound for all unbiased estimators even with the knowledge of the SST, which confirms the optimality of our proposed estimator.
While we have focused on the unbiased estimator for a specific ATE, decision-makers can have many other design choices for A/B testing in dynamic systems. For example, in practice, a slight bias may be acceptable if it substantially reduces variance. Motivated by this, we define a broad class of estimators, called differentiable estimators, which includes many existing estimators in the literature (e.g., Farias et al. 2022, 2023, Huang et al. 2023). We show that information sharing universally reduces variance without increasing bias for all differentiable estimators. Consequently, any proposed differentiable estimators, regardless of its specific design objective, could immediately benefit from the variance reduction through information sharing if the treatment is local. When there are multiple testing arms, a major component of the variance can also get a linear reduction in the number of arms. These results further illustrate the power of treatment locality.
@@
Having demonstrated the variance reduction achieved through SST, we also expand our framework and generalize the results to cover more general local treatments, which are applicable to multiple states, broadening the scope and applicability of our results. We also test the effectiveness of our design on a synthetic environment as well as a Sepsis simulator based on real-world data. Finally, we reveal that under SST, the treatment effect remains consistently positive (or negative) for all states before and after the intervention. This indicates that if the new treatment benefits one state, it benefits all others as well.
@@ 1.3 Related Works
Structure-aware Experiment Design. Experimental design originated in economics with basically model-free methods. In recent years, researchers in operations research and management science instead realize that there are many useful structures in operations scenarios to understand, analyze and improve experimentation. Bojinov et al. (2023), Hu and Wager (2022), Ni et al. (2023), Jia et al. (2023), Xiong et al. (2024) consider the switchback design under various scenarios, from temporal interference to spatial interference, from stationary environment to non-stationary environment, and from theory driven perspective to data driven perspective. Glynn et al. (2020), Munro et al. (2024), Wager and Kuang (2021), Bajari et al. (2021), Johari et al. (2022), Zhu et al. (2024), Weng et al. (2024) study the experiments happening on two-sided markets and a common practice is to model the system as a continuous-time Markov chain. Chen et al. (2025a) investigate the A/B testing in e-commerce where the product stock-out can introduce extra bias. Chen et al. (2025b) analyze the bias of experiments for inventory control policies. Li et al. (2015) leverage the sparsity structure of cross-product demand elasticities to greatly reduce the number of experiments required. Li et al. (2023) carefully take the queuing structure into account and propose efficient estimators under the stochastic congestion. Xiong et al. (2023) investigate multi-unit settings with varying treatment starting times. Zhan et al. (2021, 2023) study the post-experiment inference for contextual bandits, which does not capture the state transition as our MDPs. Simester et al. (2020) pioneer at considering treatment structures and suggest comparing outcomes only for customers exposed to different treatment and control actions to improve efficiency empirically. Compared to theirs, we further incorporate complex transition dynamics and provide theoretical guarantees of estimations. The long-term effect with the long-term treatment under the short-term measurement is recently studied by Huang et al. (2023). This paradigm of structure-aware experiment design has witnessed great success in theory and practice in above works with either significant reduction of bias or substantial improvement of efficiency.
@@
Experimenting on MDPs with General Treatments. Experimenting on MDPs is more relevant to our work. Many works adopt MDPs to capture the dynamical system and investigate the long-term outcomes (Pfeifer and Carraway 2000, Murphy 2003, Simester et al. 2006, Tran et al. 2023, Shi et al. 2020, 2023, Farias et al. 2022, 2023), where most previous works consider general treatment, i.e. a treatment will be posed to every state of MDP. The seminal work by Farias et al. (2022) studies the average reward setting. They prove the asymptotic lower bound for unbiased inference on MDPs and show that this lower bound can sometimes grow exponentially with the number of states. They also construct an LSTD(0)-type OPE estimator that is efficient and achieves this lower bound. In our work, we show that similar phenomena can also be observed in the discounted reward settings. Moreover, Farias et al. (2022) study the bias of a Naive estimator and propose an alternative estimator that achieves a notable bias-variance trade-off with an upper bound of asymptotic variance. In the sequel, we demonstrate all three estimators studied in Farias et al. (2022) fall into a general class of differentiable estimator and can thus all benefit from the variance reduction leveraging treatment locality. Finally, some recent work like Shi et al. (2023), Zhang et al. (2023) considers asymptotic results with linear function approximation. In contrast, this paper focuses on a fundamental split of the experimental data and is therefore orthogonal to previous literature on estimator design.
@@ 1.4 Structures and Notations
In Section 2, we formally introduce the model and the formulation, introducing the MDP, the treatment policy and the control policy, and the average treatment effect that we care about. Section 2.1 presents the estimation methods for the long-term cumulative rewards of a single, specific policy. In Section 3, we discuss the idea of classical A/B testing in our setting and how to estimate the long-term cumulative rewards for the treatment and control policies from A/B testing. Section 4 provides a new estimator based on the idea of information sharing for the local treatment. Section 5 generalizes the analysis to the general local treatments. Section 6 introduces the differentiable estimators, which can always benefit from the information sharing. In Section 7, we provide some numerical results verifying the idea of information sharing including the real-world sepsis simulation results. Section 8 covers the discussion on some useful properties of the local treatment as well as future work. We conclude our paper in Section 9.
@@ 2 Model
where $0\leq\gamma<1$ is a discount factor that reflects the trade-off between instant and future rewards. The value function has been widely applied to quantify long-term outcomes in various fields including marketing (Pfeifer and Carraway 2000, Simester et al. 2006), ride-hailing platforms (Shi et al. 2023), and video streaming service (Tran et al. 2023, Maystre et al. 2023). We further use ${\bm{V}}^{\pi}\in\mathbb{R}^{K}\coloneqq(V^{\pi}(s^{1}),\cdots,V^{\pi}(s^{K}))^{\top}$ to denote value functions of all states under policy $\pi$. Moreover, associated with each $\pi$, ${\bm{r}}^{\pi}\in\mathbb{R}^{K}$ is the expected reward under each state, defined as ${\bm{r}}^{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}r(s,a)\pi(a\mid s)$ for $s\in{\mathcal{S}}$. ${\bm{P}}^{\pi}\in\mathbb{R}^{K\times K}$ is the transition matrix under $\pi$ with ${\bm{P}}^{\pi}(s,s^{\prime})\coloneqq\sum_{a\in{\mathcal{A}}}P(s^{\prime}\mid s,a)\pi(a\mid s)$ for $s,s^{\prime}\in{\mathcal{S}}$.
We now introduce a special class of treatment policies that we call the Single-State Treatment (SST). Specifically, there exists a pre-determined “crucial” state, which without loss of generality, we assume to be $s^{1}$. The treatment of interest is applied exclusively when the system’s state, $s_{t}$, coincides with $s^{1}$. Formally, we assume that ${\mathcal{A}}=\{{\mathfrak{c}},{\mathfrak{t}}\}$ regardless of states, while the action $\mathfrak{t}$}, representing the new treatment, is only applicable in the crucial state $s^{1}$, while $\mathfrak{c}$} denotes the control or default action. This restriction ensures that the treatment is targeted and its effects are isolated to the crucial state.
@@
2.1 Preliminaries: Inference under a Fixed Policy
In this section, we briefly introduce how to perform inference on the value function under a fixed policy $\pi$ which is not a new task in the literature, as an important cornerstone. There are two streams of methods, both of which are reliant on the well-known Bellman Equation, i.e.,